Question 16 What is the value of x after the following code is executed? X = 7 If x < 7 X = x + 1 X = x + 2 7 9 10 8
Answer #1

Solution:-

Correct option - 9

As x is intialised to 7 and the condition is x<7 so the given condition fails and last statement x=x+2 will be executed so x = 7+2 = 9.